ashley99:

The equation the quantity x plus 33 end quantity over 4 equals 3 times x models the workload of a class project, where x is the number of hours each student must contribute. How many hours does each student work on the project? 2 hours 3 hours 7 hours 11 hours

ashley99:

Need this answer fast!!!

HawksDaughet8732:

It's either 7 or 11, bc 4+3=7 and 7+7=14-3=11

ashley99:

So, should I enter 11 or 7?

ashley99:

?